Please help me solve this word problem using the position equation. I have solved the first part but cannot seem to get the second part!
A C-141 Starlifter flying at 25,000 feet over level terrain drops a 500-pound supply package.
(a) How long will it take until the supply package strikes the ground?
This is what I have come up with:
s=-16t^2+v0t+s0
Since we don't know the velocity, plug in 0 for v0t.
s=-16t^2+0+25,000
0=-16t^2+25,000
16t^2=25,000
t^2=25,000/16
square root of t^2=square root of 25,000/16
t approximately = 39.5 seconds
Now, here's the part I can't seem to figure out.
(b) The plane is flying at 500 miles per hour. How far will the supply package travel horizontally during it's descent?

Answer by Alan3354(69443)

Please help me solve this word problem using the position equation. I have solved the first part but cannot seem to get the second part!
A C-141 Starlifter flying at 25,000 feet over level terrain drops a 500-pound supply package.
(a) How long will it take until the supply package strikes the ground?
This is what I have come up with:
s=-16t^2+v0t+s0
Since we don't know the velocity, plug in 0 for v0t.
s=-16t^2+0+25,000
0=-16t^2+25,000
16t^2=25,000
t^2=25,000/16
square root of t^2=square root of 25,000/16
t approximately = 39.5 seconds
Now, here's the part I can't seem to figure out.
(b) The plane is flying at 500 miles per hour. How far will the supply package travel horizontally during it's descent?
------------
500 mi/hr * 39.5 sec * 1 hr/3600 sec = 5.486 miles
-----------
BTW, that's called "air mail"